OpenStudy (vjohnny):

What is the Smallest Rationalizing factor of Square root of 50

OpenStudy (anonymous):

its that term when multiplied gives us rational expression in numerator or denominator

OpenStudy (unklerhaukus):

\[\sqrt{50}=\sqrt{5^2\times2}=5\sqrt2\]

OpenStudy (vjohnny):

Thanks, So \[\sqrt{2}\] would be the Rationalizing factor, since multiplying \[\sqrt{50}\] by \[\sqrt{2}\] would give us a Rational number which is 10